OpenStudy (anonymous):

1. Which of the following is a factor of x^2 – 6x – 27? a. x + 3 b. x + 9 c. x – 1 d. None of the above

OpenStudy (across):

Have you tried synthetic division?

OpenStudy (jiteshmeghwal9):

use the identity:-\[x^2-(a+b)x+ab\]

OpenStudy (anonymous):

@across, no.

OpenStudy (anonymous):

I would use REMAINDER THEOREM

OpenStudy (anonymous):

Do u know it? @bombshellbri

OpenStudy (anonymous):

never learned that...

OpenStudy (anonymous):

substitute the given factors ....

OpenStudy (anonymous):

Actually Factor Theorem

OpenStudy (anonymous):

If polynomial f(x) is divided by (x-a) then (x-a) is said to be a factor of polynomial f(x) if f(a)=0

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

not really.

OpenStudy (anonymous):

Here: f(x)=x^2-6x-27

OpenStudy (anonymous):

no i mean to say that suppose to test x-a substitute x=a in the given polynomial and if reduces to zero then x-a is its factor

OpenStudy (anonymous):

@sauravshakya , so where f(x) is, i should substitute the given choices to replace x?

OpenStudy (anonymous):

(x-a) ===> (x+3) a=-3

OpenStudy (anonymous):

example to for x^2-6x-27 testing x+3 let x+3=0 then x=-3 substitute that -3^2-6*(-3)-27 it gives 0 hence its a factor

OpenStudy (anonymous):

Now, find f(a)

OpenStudy (anonymous):

x + 3 is a factor

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

so this is the answer. correct?

OpenStudy (anonymous):

Yes very correct

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

Welcome